Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

String of lightbulbs

Expert replies
by jfranco23 » Tue Mar 17, 2009 5:06 pm
A string of lightbulbs is wired in such a way that if any individual lightbulb fails, the entire string fails. If for each individual lightbulb the probability of failing during time period T is 0.06, what is the probability that the string of lightbulbs will fail during time period T?

A. 0.06
B. (0.06)^10
C. 1 - (0.06)^10
D. (0.94)^10
E. 1 - (0.94)^10
Join the discussion
Source: — Problem Solving |

by cubicle_bound_misfit » Wed Mar 18, 2009 7:52 am
whats the no of lightbulbs in the string?
Cubicle Bound Misfit
Join the discussion

by rseeker2 » Wed Mar 18, 2009 8:49 am
probability that the string of lightbulbs will fail during time period T --> 1-P(no bulb failing)

= 1 - (0.94)^n

If n were 10, it would be 1 - (0.94)^10
Join the discussion

by Stuart@KaplanGMAT » Wed Mar 18, 2009 11:00 am
rseeker2 wrote:probability that the string of lightbulbs will fail during time period T --> 1-P(no bulb failing)

= 1 - (0.94)^n

If n were 10, it would be 1 - (0.94)^10
Perfect!

Based on the answer choices, there must be 10 lightbulbs - the OP must have missed part of the question (maybe it really started with "A string of 10 lightbulbs...").
Image

Stuart Kovinsky | Kaplan GMAT Faculty | Toronto

Kaplan Exclusive: The Official Test Day Experience | Ready to Take a Free Practice Test? | Kaplan/Beat the GMAT Member Discount
BTG100 for $100 off a full course
Join the discussion

by jfranco23 » Wed Mar 18, 2009 11:58 am
So sorry for my mistake, there are 10 lightbulbs.
Join the discussion

by hijazim » Thu Mar 19, 2009 7:09 am
Shouldn't it be (0.06)x10 because it is OR not AND probability. As if any of the lightbulbs failed the whole string will fail, so if the 1st or 2nd or ... 10th fail the wire will fail..

so it has to be : 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 = 0.6

please advise.
Join the discussion

by sureshbala » Thu Mar 19, 2009 7:42 am
hijazim wrote:Shouldn't it be (0.06)x10 because it is OR not AND probability. As if any of the lightbulbs failed the whole string will fail, so if the 1st or 2nd or ... 10th fail the wire will fail..

so it has to be : 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 + 0.06 = 0.6

please advise.
Folk in addition to this the light will fail if 1st AND 2nd fail and the other don't fail. 1st, 2nd AND 3rd fail and the rest don't....so on...

So instead of considering all this we calculate the probability that none of the light fails for which the probability is (0.94)^10.

Hence the probability that at least one of them fails = 1- (0.94)^10
Join the discussion